OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? −x + 2y = 9 4x + 5y = 9 −9 + 2y −9 − 2y 9 − 2y 9 + 2y

jimthompson5910 (jim_thompson5910):

if you solved −x + 2y = 9 for x, what would you get?

OpenStudy (anonymous):

i dont know

jimthompson5910 (jim_thompson5910):

I'll start you off −x + 2y = 9 −x + 2y - 2y = 9 - 2y ... Subtract 2y from both sides. -x = 9 - 2y and I'll let you finish up

jimthompson5910 (jim_thompson5910):

you need to completely isolate x

OpenStudy (anonymous):

9-2y

jimthompson5910 (jim_thompson5910):

that's -x NOT x itself

OpenStudy (anonymous):

so how do you get rid of the -x and make x itself

jimthompson5910 (jim_thompson5910):

just multiply everything by -1

jimthompson5910 (jim_thompson5910):

what do you get when you do this?

OpenStudy (anonymous):

-9+ 2y

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

so x = -9 + 2y

OpenStudy (anonymous):

thank you :]

jimthompson5910 (jim_thompson5910):

which means that you would substitute that into the second equation

OpenStudy (anonymous):

would you mind helping me with this next one ?

jimthompson5910 (jim_thompson5910):

sure i can help

OpenStudy (anonymous):

Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 4x + 3y = -1 3x + y = 3

jimthompson5910 (jim_thompson5910):

substitution would be easiest, so let's use that

jimthompson5910 (jim_thompson5910):

Solve equation 2 for y 3x + y = 3 3x + y - 3x = 3 - 3x y = 3 - 3x --------------------------- 4x + 3y = -1 4x + 3(3 - 3x) = -1 ... plug in y = 3 - 3x 4x + 9 - 9x = -1 keep going to solve for x

jimthompson5910 (jim_thompson5910):

what do you get